OpenStudy (queenof17):

If the equation of a circle is (x + 4) 2 + (y - 6) 2 = 25, its radius is

OpenStudy (priyar):

its in the standars format of : (x-h)^2 + (y-k)^2 = r^2 where (h,k)-centre r- radius So what will be the answer?

OpenStudy (studio-games):

x=−y+ 29/2

OpenStudy (studio-games):

hope that helped.

OpenStudy (queenof17):

Could you walk me through how to get the answer? Everything about geometry confuses me... @studio-games @priyar

OpenStudy (studio-games):

Let's solve for x. (x+4)(2)+(y−6)(2)=25 Step 1: Add 4 to both sides. 2x+2y−4+4=25+4 2x+2y=29 Step 2: Add -2y to both sides. 2x+2y+−2y=29+−2y 2x=−2y+29 Step 3: Divide both sides by 2.

OpenStudy (studio-games):

2x/2 = −2y+29/2 x= −y + 29/2

OpenStudy (studio-games):

There you go.

OpenStudy (queenof17):

My answers could either be one of the following: 5, 10, or 25

OpenStudy (queenof17):

Does that change the way I need to solve the problem?

OpenStudy (freckles):

@queenof17 do you know the distance formula If I wrote \[5=\sqrt{(x-(-4))^2+(y-6)^2}\] could you tell me the distance between (x,y) and (-4,6) ?

OpenStudy (queenof17):

Could you explain how I would go about solving it?

OpenStudy (queenof17):

@freckles

OpenStudy (freckles):

do you know the distance formula?

OpenStudy (freckles):

\[\text{The distance from } \\ (x_1,y_1) \text{ to } (x_2,y_2) \\ \text{ is given by } d \\ \text{ in } \\ d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (freckles):

Any point (x,y) on a circle has the same distance from the center of the circle This distance is known as the radius

OpenStudy (queenof17):

Okay, so what would I do first to solve it?

OpenStudy (freckles):

we have \[5=\sqrt{(x-(-4))^2+(y-6)^2} \\ \text{ looking at this } \\ \text{ can you tell me the distance from } \\ (x,y) \text{ to } (-4,6)\]

OpenStudy (freckles):

compare this to the distance formula I wrote above

OpenStudy (queenof17):

Was I supposed to get 20 and 36

OpenStudy (freckles):

are you comparing \[5=\sqrt{(x-(-4))^2+(y-6)^2} \\ \text{ to } \\ d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] I'm asking you for the distance between (x,y) and (-4,6) \[\text{ the first equation says the distance between } (x,y) \text{ and } (-4,6) \text{ is ....}\] \[\text{ the second equation says the distance between } (x_1,y_1) \text{ and } (x_2,y_2) \text{ is } d \]

OpenStudy (freckles):

what is in place of d in the first equation?

OpenStudy (queenof17):

5,

OpenStudy (freckles):

yes!

OpenStudy (mathmale):

Afraid we're going into far more detail here than is ncessary! You are given " (x + 4) 2 + (y - 6) 2 = 25 " and can / must rewrite this as \[(x+4)^2+(y-6)^2=25=5^2\]

OpenStudy (freckles):

the distance between any point (x,y) and (-4,6) is 5

OpenStudy (mathmale):

compare this to the general equation of a circle:\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (mathmale):

Simply by doing so, you can immediately see what the value of r is. Try it. (h,k) represents the center of the circle, by the way.

OpenStudy (freckles):

@mathmale I think @priyar tried that approach of just comparing to (x-h)^2+(y-k)^2=r^2 that is why I tried something else

OpenStudy (queenof17):

I just need a step by step on how to solve the problem. I'm awful at geometry and I have a lot of problems like this

OpenStudy (mathmale):

If you want to learn a different way of solving this problem, go ahead. I don't see any particular advantage in using the distance formula here, but it is certainly applicable. Again I think you're making this problem overly complicated.

OpenStudy (mathmale):

I just gave you a "step by step" approach. Compare YOUR equation with the STANDARD equation; you can read off the radius, r, immediately.

OpenStudy (queenof17):

I am, I just can't see how the number 25 is brought in

OpenStudy (queenof17):

Again, I'm not great at geometry

OpenStudy (mathmale):

Why do more than that if it's not necessary? You are GIVEN that 25. That 25 is the SQUARE of the radius. Thus, the radius is 5. End of discussion. Kindly STOP that "I'm not great at geometry" stuff. What do you hope to accomplish by saying that? Give yourself credit for your intelligence and ability to learn.

OpenStudy (mathmale):

Please, move on to another problem. Post it separately from this one.